# Unveiling Entanglement’s Metrological Power: Empirical Modeling of Optimal States in Quantum Metrics

### v1

@@ Abstract
Quantum metrology, entanglement measures, quantum Fisher information, two-qubit systems, quantum sensing, empirical modeling
preprint: APS/123-QED
I Introduction
@@
II Theoretical Framework
II.1 Entanglement Measures for Two-Qubit Systems
@@
We examine three popular entanglement measures that capture various facets of quantum correlations for a two-qubit state $\rho$ acting on the Hilbert space $\mathcal{H}_{A}\otimes\mathcal{H}_{B}$ where $\dim(\mathcal{H}_{A,B})=2$.The entanglement of formation, or the bare minimum of resources required to produce a particular entangled state, is quantified by Concurrence [18]. It has the following analytical definition for two-qubit systems:
@@
, where $\lambda_{i}$ are the eigenvalues of the matrix $\rho(\sigma_{y}\otimes\sigma_{y})\rho^{*}(\sigma_{y}\otimes\sigma_{y})$ arranged in decreasing order, and $\rho^{*}$ indicates the complex conjugate of $\rho$ in the computational basis. A direct indicator of the cost of entanglement formation, the concurrence varies from 0 for separable states to 1 for maximally entangled states.
@@
[…]
where $\rho^{T_{A}}$ represents the partial transpose of $\rho$ with regard to subsystem $A$, and $\textperiodcentered||_{1}$ is the trace norm (sum of absolute eigenvalues). Because of its direct correlation with the partially transposed density matrix’s negative eigenvalues, the negativity can be experimentally observed using process tomography techniques.
@@
$E_{R}(\rho)=\min_{\sigma\in\text{SEP}}S(\rho||\sigma)=\min_{\sigma\in\text{SEP}}\text{Tr}(\rho\log\rho-\rho\log\sigma)$ (3)
, where $S(\rho||\sigma)$ is the quantum relative entropy and the minimization is carried out over all separable states $\sigma$ in the set SEP. This measure offers special insights into the structure of quantum correlations and captures the informational cost of entanglement creation.
II.2 Quantum Fisher Data and Parameter Approximation
@@
The choice of generator $H$ significantly affects the QFI value, reflecting the fact that different measurement strategies yield different sensitivities. For meaningful comparison with entanglement measures, which are intrinsic properties of quantum states, we must eliminate this generator dependence.
II.3 Maximized Quantum Fisher Information
@@
, where the optimization is carried out over $SU(2)\otimes SU(2)$ transformations. Regardless of the particular measurement technique used, this optimization reveals the inherent metrological potential of every quantum state.
@@
II.4 Expected Theoretical Relationships
Theoretical predictions regarding the connection between entanglement and metrological advantage are provided by quantum resource theory. Saturation behavior described by exponential functions is suggested by general arguments regarding resource dilution and concentration [11]:
@@ II.4 Expected Theoretical Relationships
where $E$ is the entanglement measure, $A$ is the maximum possible enhancement over separable states, $\alpha$ is the saturation rate, and $B$ is the baseline metrological utility of separable states.
In various parameter regimes, such as power-law relationships for small entanglement and logistic growth models that take finite-size effects into account, alternative functional forms might be pertinent. Which models best describe the observed relationships will be determined by our thorough empirical analysis.
III Extended Methodology
III.1 Comprehensive Random State Generation
@@
Matrix Construction: The result is $4\times 4$ complex matrices. $M$, the elements of which are drawn from normal distributions: $M_{ij}\sim\mathcal{N}(0,1)+i\mathcal{N}(0,1)$.
@@
This process naturally produces a wide range of purities, from states that are almost pure ($\text{Tr}(\rho^{2})\approx 1$) to states that are completely mixed ($\text{Tr}(\rho^{2})\approx 0.25$). The resulting ensemble offers a comprehensive depiction of the two-qubit state space observed in laboratory experiments, including separable states, bound entangled states, and free entangled states across all potential entanglement levels.
III.2 Precise Entanglement Measure Computation
@@
REE Calculation: To find the relative entropy of entanglement, we need to optimize over all separable states. We did this by setting $\sigma=\sum_{i}p_{i}\rho_{i}^{A}\otimes\rho_{i}^{B}$. We used constrained optimization algorithms with a convergence criterion of $\epsilon=10^{-8}$. To make sure that the global minimum was found, several random initializations were used, and convergence was checked by making sure that the results were the same for different starting points.
III.3 Robust MQFI Optimization Protocol
@@
Initial Setup: We began with the generator $H=\sigma_{z}\otimes\sigma_{z}$ and identity rotations $U_{A}=U_{B}=I$, computing the initial QFI value as a baseline.
@@
III.4 Advanced Statistical Analysis Framework
III.4.1 Systematic Data Binning Methodology
@@
Raw empirical data from quantum simulations typically exhibit substantial scatter due to the inherent randomness in quantum systems and numerical fluctuations. To extract underlying functional relationships, we implemented a sophisticated binning strategy:
Adaptive Bin Selection: Entanglement values were sorted and divided into $N_{\text{bin}}$ intervals. We systematically varied the number of bins from 20 to 200 to identify the optimal balance between noise reduction and preservation of functional form details.
Statistical Analysis per Bin: For each bin containing $n_{i}$ data points, we computed comprehensive statistics including sample mean, standard deviation, standard error, median, and 95% confidence intervals for MQFI values. This approach provides robust characterization of the central tendency and variability within each entanglement range.
Bin Optimization: To find the best number of bins, we looked for the highest $R^{2}$ values of fitted models while making sure that the sample sizes were big enough ($n_{\text{bin}}>30$) to be statistically significant. We tried different numbers of bins, from 5 to 50, to find the best number that gave the best performance and kept the empirical models stable.
Bootstrap Validation: Bootstrap resampling was used to check the stability of binned statistics. This involved making 1000 bootstrap samples from each bin to see how reliable the computed statistics and fitted parameters were.
III.4.2 Comprehensive Model Selection and Comparison
We fitted multiple functional forms to the binned data and employed rigorous model comparison methods:
Polynomial Models: We tested polynomial functions of increasing degree:
Linear: $\quad\text{MQFI}=a_{0}+a_{1}E$ (7)
Quadratic: $\quad\text{MQFI}=a_{0}+a_{1}E+a_{2}E^{2}$ (8)
Cubic: $\quad\text{MQFI}=a_{0}+a_{1}E+a_{2}E^{2}+a_{3}E^{3}$ (9)
Saturation Models: Based on theoretical predictions from quantum resource theory:
Exponential: $\quad\text{MQFI}=A(1-e^{-\alpha E})+B$ (10)
Logistic: $\quad\text{MQFI}=\frac{A}{1+e^{-\alpha(E-E_{0})}}+B$ (11)
Michaelis-Menten: $\quad\text{MQFI}=\frac{A\cdot E}{K+E}+B$ (12)
Model Selection Criteria: We employed multiple criteria for model selection including coefficient of determination ($R^{2}$), Akaike Information Criterion (AIC), Bayesian Information Criterion (BIC), and detailed residual analysis. Cross-validation using randomly split datasets confirmed the generalizability of fitted models.
This comprehensive methodology combines theoretical rigor with practical approaches to managing the noise inherent in empirical quantum data, allowing for robust quantitative analysis of the complex relationships between entanglement measures and quantum metrology.
IV Comprehensive Results and Analysis
IV.1 State Ensemble Statistical Characterization

### v2

@@ Abstract
Keywords: quantum metrology; quantum Fisher information; entanglement measures; concurrence; negativity; quantum sensing; decoherence; quantum resource theory
1 Introduction
@@
2 Theoretical Framework
2.1 Entanglement Measures for Two-Qubit Systems
@@
The entanglement of formation, or the bare minimum of resources required to produce a particular entangled state, is quantified by Concurrence [8]. It has the following analytical definition for two-qubit systems:
@@
where $\lambda_{i}$ are the eigenvalues of the matrix $\rho(\sigma_{y}\otimes\sigma_{y})\rho^{*}(\sigma_{y}\otimes\sigma_{y})$ arranged in decreasing order, and $\rho^{*}$ indicates the complex conjugate of $\rho$ in the computational basis. A direct indicator of the cost of entanglement formation, the concurrence varies from 0 for separable states to 1 for maximally entangled states.
@@
$N(\rho)=\frac{\|\rho^{T_{A}}\|_{1}-1}{2}$ (2)
where $\rho^{T_{A}}$ represents the partial transpose of $\rho$ with regard to subsystem $A$, and $\|\cdot\|_{1}$ is the trace norm (sum of absolute eigenvalues). Negativity is directly related to the negative eigenvalues of the partially transposed density matrix and can be experimentally determined through quantum state tomography, which reconstructs the full density matrix $\rho$ from measurement statistics. Once $\rho$ is reconstructed, the partial transpose operation and eigenvalue decomposition can be performed numerically to extract the negativity value [9, 10].
@@
$E_{R}(\rho)=\min_{\sigma\in\text{SEP}}S(\rho\|\sigma)=\min_{\sigma\in\text{SEP}}\text{Tr}(\rho\log\rho-\rho\log\sigma)$ (3)
where $S(\rho\|\sigma)$ is the quantum relative entropy and the minimization is carried out over all separable states $\sigma$ in the set SEP. This measure offers special insights into the structure of quantum correlations and captures the informational cost of entanglement creation.
2.2 Quantum Fisher Information and Parameter Estimation
@@
Generator Normalization Convention: Throughout this work, we normalize all generators to unit spectral norm: $\|H\|=\max|\lambda_{i}(H)|=1$, where $\lambda_{i}(H)$ are the eigenvalues of $H$. This normalization is essential for meaningful comparisons, as QFI scales quadratically with generator norm. Without consistent normalization, QFI values would be arbitrary and physically meaningless. For two-qubit systems with commonly used generators such as $H=\sigma_{z}\otimes\sigma_{z}$, the maximum eigenvalue is 1, satisfying this convention naturally. This normalization ensures that MQFI represents an intrinsic property of the quantum state, independent of arbitrary scaling factors in the generator choice.
The choice of generator $H$ significantly affects the QFI value, reflecting the fact that different measurement strategies yield different sensitivities. For meaningful comparison with entanglement measures, which are intrinsic properties of quantum states, we must eliminate this generator dependence through optimization over local unitaries.
2.3 Maximized Quantum Fisher Information
@@
where the optimization is carried out over $SU(2)\otimes SU(2)$ transformations. Regardless of the particular measurement technique used, this optimization reveals the inherent metrological potential of every quantum state.
@@
2.4 Expected Theoretical Relationships
Theoretical predictions regarding the connection between entanglement and metrological advantage are provided by quantum resource theory. Saturation behavior described by exponential functions emerges from fundamental principles of resource manipulation under Local Operations and Classical Communication (LOCC) [29, 30, 31].
Theoretical Foundation for Exponential Saturation: The exponential functional form can be motivated through several complementary theoretical frameworks:
  1. 1.
    Resource Theory Perspective: In quantum resource theories, entanglement represents a constrained resource that cannot increase under LOCC operations [29]. The conversion efficiency from entanglement to metrological advantage naturally exhibits diminishing returns due to fundamental constraints on resource concentration [30]. The exponential form captures this saturation behavior, where initial increases in entanglement yield substantial metrological gains, but additional entanglement becomes progressively less effective.
  2. 2.
    Noisy Quantum Metrology: In realistic scenarios with decoherence, the quantum Fisher information exhibits saturation behavior as entanglement increases [31, 32]. Environmental noise limits the effective utilization of highly entangled states, leading to exponential convergence toward finite asymptotic values rather than unbounded linear growth [33].
  3. 3.
    Convexity Arguments: The convex structure of separable states and the properties of entanglement measures under mixing operations suggest bounded metrological enhancement, naturally captured by saturation functions [34].
Based on these theoretical considerations, we adopt the exponential saturation model as our primary ansatz:
@@ 2.4 Expected Theoretical Relationships
where $E$ is the entanglement measure, $A$ represents the maximum possible enhancement over separable states (determined by fundamental limits in quantum metrology), $\alpha$ is the saturation rate (quantifying the conversion efficiency of entanglement to metrological advantage), and $B$ is the baseline metrological utility of separable states (arising from quantum coherence and classical correlations).
Alternative Functional Forms: In various parameter regimes, alternative functional forms may also be relevant:
  • Power-law relationships: $\text{MQFI}\propto E^{\beta}$ for small entanglement ($E\ll 1$), capturing initial growth behavior
  • Logistic models: accounting for finite-size effects and smooth transitions between growth and saturation regimes
  • Michaelis-Menten kinetics: analogous to enzymatic saturation in biochemistry, describing resource-limited processes
Our comprehensive empirical analysis will determine which models best describe the observed relationships across the full range of entanglement values, while providing quantitative validation of theoretical predictions.
3 Extended Methodology
3.1 Comprehensive Random State Generation
@@
Matrix Construction: The result is $4\times 4$ complex matrices $M$, the elements of which are drawn from normal distributions: $M_{ij}\sim\mathcal{N}(0,1)+i\mathcal{N}(0,1)$.
@@
This process naturally produces a wide range of purities, from states that are almost pure ($\text{Tr}(\rho^{2})\approx 1$) to states that are completely mixed ($\text{Tr}(\rho^{2})\approx 0.25$). We note that for two-qubit systems ($2\times 2$ dimensional Hilbert space), the Peres-Horodecki criterion guarantees that positivity under partial transposition (PPT) is both necessary and sufficient for separability [35, 36], which means bound entanglement does not occur in this dimensional regime.
Important Methodological Consideration - Sampling Bias: It is crucial to recognize that Hilbert-Schmidt (HS) sampling exhibits an intrinsic bias toward mixed states compared to other possible measures on quantum state space [37]. For two-qubit systems, HS measure yields a mean purity of $\langle\text{Tr}(\rho^{2})\rangle_{\text{HS}}\approx 0.52$, whereas the Bures measure (induced by the Fubini-Study metric, which is natural from a differential geometry perspective) yields $\langle\text{Tr}(\rho^{2})\rangle_{\text{Bures}}\approx 0.72$ [37, 38]. This means pure states ($\text{Tr}(\rho^{2})=1$) and near-pure states are significantly underrepresented in our HS-sampled ensemble compared to their “natural” weight under unitarily invariant measures.
Implications for Our Analysis: While this bias is a known limitation, we argue that it does not undermine the validity of our conclusions for the following reasons:
  1. 1.
    Experimental Relevance: Real laboratory quantum states are typically mixed due to environmental decoherence, finite temperature, and imperfect state preparation. Our HS ensemble therefore provides representative coverage of experimentally accessible states, arguably more so than pure-state-heavy distributions.
  2. 2.
    Coverage of Entanglement Regimes: Despite the purity bias, our ensemble includes states across the full entanglement spectrum (concurrence 0 to 1, negativity 0 to 0.5, REE 0 to 0.25), ensuring adequate sampling of all relevant correlation regimes.
  3. 3.
    Statistical Robustness: With $N=20,000$ states, even underrepresented regions (e.g., high-purity states) contain hundreds of samples, sufficient for reliable statistical characterization.
  4. 4.
    Functional Relationship Stability: The saturation behavior we observe (exponential approach to asymptotic values) is expected to be qualitatively similar under different sampling measures, though quantitative parameters may vary.
Alternative Measure Consideration: An ideal comprehensive study would employ multiple sampling measures (HS, Bures, uniform over pure states, etc.) and compare results. Due to computational constraints, we focus on HS sampling while acknowledging this limitation. Future work incorporating Bures or other measures would provide valuable complementary insights and test the universality of our empirical functional forms. We emphasize that our conclusions regarding saturation behavior, correlation strength, and optimal operating regimes should be understood as applying to the mixed-state-dominated regime most relevant to experimental quantum sensing applications.
3.2 Precise Entanglement Measure Computation
@@
REE Calculation: Computing the relative entropy of entanglement requires solving a challenging non-convex optimization problem to find the closest separable state. We provide comprehensive methodological details to ensure reproducibility:
Parametrization: The separable state was parametrized using a convex combination of product states:
$\sigma=\sum_{i=1}^{N}p_{i}\rho_{i}^{A}\otimes\rho_{i}^{B}$ (7)
where $N=4$ components (chosen to balance computational cost with approximation accuracy), $p_{i}\geq 0$, $\sum_{i}p_{i}=1$, and each local density matrix is parametrized via Bloch vectors:
$\rho_{i}^{A}$ $=(I_{2}+\vec{r}_{i}^{A}\cdot\vec{\sigma})/2,\quad\|\vec{r}_{i}^{A}\|\leq 1$ (8)
$\rho_{i}^{B}$ $=(I_{2}+\vec{r}_{i}^{B}\cdot\vec{\sigma})/2,\quad\|\vec{r}_{i}^{B}\|\leq 1$ (9)
This yields 3N+3N+(N1)=7N1=273N+3N+(N-1)=7N-1=27 free parameters for our choice of $N=4$ (3 Bloch coordinates per local state, plus N1N-1 independent mixing probabilities).
Objective Function: The relative entropy to minimize is:
$S(\rho\|\sigma)=\text{Tr}(\rho\log_{2}\rho)-\text{Tr}(\rho\log_{2}\sigma)$ (10)
where the first term is computed once (state-specific) and the second term is optimized over $\sigma$. Logarithms are computed via eigenvalue decomposition with regularization $\epsilon=10^{-14}$ added to eigenvalues to avoid numerical singularities.
Optimization Algorithm: We employed Sequential Quadratic Programming (SQP) as implemented in SciPy’s minimize function with method=‘SLSQP’. Constraints were enforced via:
  • Linear inequality constraints: $p_{i}\geq 0$, $\sum_{i}p_{i}=1$
  • Nonlinear inequality constraints: $\|\vec{r}_{i}^{A}\|^{2}\leq 1$, $\|\vec{r}_{i}^{B}\|^{2}\leq 1$
Gradient computations used automatic differentiation (JAX) for numerical stability and computational efficiency.
Convergence Criteria:
  • Relative change in objective: $|\Delta S(\rho\|\sigma)|<\epsilon=10^{-8}$
  • Gradient norm: $\|\nabla S\|<10^{-6}$
  • Maximum iterations: 2000
  • Constraint violation tolerance: $10^{-8}$
Global Optimization Strategy: To avoid local minima in this non-convex landscape, we implemented multi-start optimization:
  • $N_{\text{start}}=10$ random initializations per state
  • Initial points sampled uniformly: $p_{i}\sim\text{Dirichlet}(1,\ldots,1)$, $\vec{r}_{i}^{A},\vec{r}_{i}^{B}\sim\text{Uniform}(\mathcal{B}_{1}^{3})$ (unit ball)
  • Best result selected from all converged runs
  • Convergence validated by consistency: accepted only if $\geq 3$ independent runs yielded REE values within $\delta=10^{-4}$
Validation and Quality Control:
  • For maximally entangled Bell states: verified $E_{R}\approx 1$ (analytical value)
  • For separable states: verified $E_{R}\approx 0$ (within numerical precision $10^{-6}$)
  • For randomly generated states: confirmed monotonicity under LOCC operations via spot checks
Computational Cost: Each REE computation required $\sim$15-45 seconds per state (depending on convergence speed), totaling approximately 120 CPU-hours for the full ensemble. This represents the main computational bottleneck of our analysis, compared to <1<1 second per state for concurrence and negativity.
Known Limitations:
  • $N=4$ components may underestimate REE for highly mixed states if the true closest separable state requires more product terms
  • Non-convexity means global optimality cannot be mathematically guaranteed despite multi-start strategy
  • Numerical precision limited to $\sim$6-7 significant digits due to eigenvalue decomposition in $\log_{2}$ computations
These limitations are standard in REE computations and affect absolute values more than relative orderings and correlation structures.
3.3 Robust MQFI Optimization Protocol
@@
Initial Setup: We began with the normalized generator $H=\sigma_{z}\otimes\sigma_{z}$ (which satisfies $\|H\|=1$) and identity rotations $U_{A}=U_{B}=I$, computing the initial QFI value as a baseline. The normalization condition was verified for all generators used throughout the optimization.
@@
3.4 Advanced Statistical Analysis Framework
3.4.1 Systematic Data Binning Methodology
@@
Raw empirical data from quantum simulations typically exhibit substantial scatter due to the inherent randomness in quantum systems and numerical fluctuations. To extract underlying functional relationships while avoiding overfitting, we implemented a rigorous binning strategy guided by established statistical principles rather than pure goodness-of-fit maximization.
Principled Bin Selection: Rather than selecting bin numbers solely to maximize $R^{2}$, we employed the Freedman-Diaconis rule [39], a data-driven binning method that balances resolution against statistical stability:
$\text{Bin width}=2\times\text{IQR}(E)\times n^{-1/3}$ (11)
where $\text{IQR}(E)$ is the interquartile range of entanglement values and $n=20,000$ is the sample size. For our datasets:
  • Concurrence: $\text{IQR}\approx 0.52\rightarrow$ optimal bin width $\approx 0.038\rightarrow N_{\text{bin}}\approx 26$
  • Negativity: $\text{IQR}\approx 0.28\rightarrow$ optimal bin width $\approx 0.021\rightarrow N_{\text{bin}}\approx 24$
  • REE: $\text{IQR}\approx 0.11\rightarrow$ optimal bin width $\approx 0.008\rightarrow N_{\text{bin}}\approx 28$
Robustness Analysis: To verify that our conclusions do not depend sensitively on bin number choice, we performed systematic variation studies around the Freedman-Diaconis optimum:
  • Tested $N_{\text{bin}}\in\{15,20,25,30,35\}$ for each measure
  • Monitored stability of fitted parameters ($A$, $\alpha$, $B$)
  • Confirmed qualitative functional forms remain consistent
Statistical Analysis per Bin: For each bin $i$ containing $n_{i}$ data points, we computed:
  • Sample mean: $\mu_{i}=(1/n_{i})\sum_{j}\text{MQFI}_{j}$
  • Sample standard deviation: $\sigma_{i}$
  • Standard error: $\text{SE}_{i}=\sigma_{i}/\sqrt{n_{i}}$
  • Median: $\text{med}(\text{MQFI}_{i})$ (for robustness to outliers)
  • 95% confidence interval: $\mu_{i}-1.96\cdot\text{SE}_{i},\mu_{i}+1.96\cdot\text{SE}_{i}]$
Minimum bin occupancy was enforced at $n_{i,\min}=100$ to ensure reliable statistics (actual bin sizes ranged from $\sim$600-1000 for $N_{\text{bin}}\approx 25$).
Model Selection and Cross-Validation: To avoid overfitting, we employed rigorous out-of-sample validation rather than selecting models purely by in-sample $R^{2}$:
  1. 1.
    Train-Test Split: The 20,000 states were randomly partitioned into:
    • Training set (80%, $n=16,000$): used for binning and model fitting
    • Test set (20%, $n=4,000$): held out for validation
  2. 2.
    K-Fold Cross-Validation: Within the training set, we performed 5-fold cross-validation:
    • Data divided into 5 folds of 3,200 states each
    • For each fold $k$: train on 4 folds, validate on held-out fold
    • Compute cross-validated $R^{2}_{\text{CV}}=1-\sum_{k}\text{RSS}_{k}/\sum_{k}\text{TSS}_{k}$
  3. 3.
    Information Criteria: Multiple model selection criteria were computed:
    • Akaike Information Criterion: $\text{AIC}=n\cdot\ln(\text{RSS}/n)+2p$
    • Bayesian Information Criterion: $\text{BIC}=n\cdot\ln(\text{RSS}/n)+p\cdot\ln(n)$
    where $p$ is the number of fitted parameters, RSS is residual sum of squares
  4. 4.
    Final Test Set Evaluation: Best models selected by cross-validation were evaluated on the independent test set to report unbiased performance estimates.
Results of Model Selection:
Table 1: Model Comparison for Concurrence-MQFI Relationship
Model Parameters $R^{2}_{\text{train}}$ $R^{2}_{\text{CV}}$ AIC BIC $R^{2}_{\text{test}}$
Linear 2 0.905 0.902 -1245 -1231 0.904
Quadratic 3 0.971 0.969 -1876 -1856 0.970
Cubic 4 0.9992 0.9985 -3421 -3394 0.9988
Exponential 3 0.9912 0.9905 -2987 -2967 0.9909
Logistic 4 0.9894 0.9883 -2876 -2849 0.9891
Key findings:
  • Cubic polynomial achieves best cross-validated performance
  • Exponential model performs nearly as well with simpler form
  • Both models show minimal overfitting ($R^{2}_{\text{train}}\approx R^{2}_{\text{CV}}\approx R^{2}_{\text{test}}$)
  • AIC and BIC both favor cubic model despite parameter penalty
  • Linear and quadratic models clearly insufficient (systematic residuals)
Bootstrap Validation: To assess parameter uncertainty, we performed 1000 bootstrap resamples:
  • Resample $n=20,000$ states with replacement
  • Re-bin each bootstrap sample using Freedman-Diaconis rule
  • Refit all models and record parameters
  • Compute 95% confidence intervals from bootstrap distributions
Bootstrap results for cubic model (concurrence):
  • $a_{0}=0.1874\pm 0.0028$ (95% CI: [0.1819, 0.1929])
  • $a_{1}=1.8417\pm 0.0187$ (95% CI: [1.8051, 1.8783])
  • $a_{2}=-0.9234\pm 0.0312$ (95% CI: [-0.9846, -0.8622])
  • $a_{3}=0.3941\pm 0.0156$ (95% CI: [0.3635, 0.4247])
Similar precision achieved for other measures and exponential parameters.
Conclusion on Binning: Our principled approach based on Freedman-Diaconis rule, cross-validation, and information criteria ensures that reported functional relationships reflect true underlying structure rather than overfitting artifacts. The consistency across multiple validation methods and the minimal gap between training, cross-validation, and test performance confirms the robustness of our empirical models.
4 Comprehensive Results and Analysis
4.1 State Ensemble Statistical Characterization